OpenStudy (anonymous):

Write the following polar equation in Cartesian coordinates: 2r(sin θ + cos θ) = 5. b) find the y-intercept and slope

OpenStudy (tkhunny):

What have you tried? Do you know the conversions? This one jumps right out at you if you do.

OpenStudy (anonymous):

I divided both sides by 2, and got the y-intercept to be 5/2. I'm not sure of the conversion so I'm a bit confused..

OpenStudy (tkhunny):

WAY too hard. \(x = r\cos(\theta)\), right?

OpenStudy (anonymous):

I also have another with Write the following polar equation in Cartesian coordinates: r = sin (2θ) but i'm stuck

OpenStudy (anonymous):

Yes x= r cos theta and y = r sin theta

OpenStudy (tkhunny):

One at a time, please. Okay, now use your Distributive Property to get rid of the parentheses.

OpenStudy (anonymous):

r sin (θ) + r cos (θ) = 5/2

OpenStudy (anonymous):

so that should be equivalent to x + y = 5/2?

OpenStudy (tkhunny):

This was just waiting for you not to see it. Now, answer the question about as fast as youcan write it, right? Slope = y-Intercept =

OpenStudy (anonymous):

slope = -1 and y-intercept = 5/2

OpenStudy (anonymous):

how will I write the cartesian coordinate?

OpenStudy (tkhunny):

Good work. Repost the other one and tag me. We can talk there.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Write the following polar equation in Cartesian coordinates: r = sin (2θ) becomes

OpenStudy (tkhunny):

I do not see exactly how to proceed. My recommendation under that circumstance is to give something a shot and see where it goes. We have some fundamental clues. \(x = r\cos(\theta)\) \(y = r\sin(\theta)\) One more thing comes to mind that might be helpful. \(\sin(2\theta) = 2\sin(\theta)\cos(\theta)\) How can we get all that together?

OpenStudy (anonymous):

x= sin 2θ (cos θ) y=sin 2θ (sin θ)

OpenStudy (tkhunny):

?? I don't see that.

OpenStudy (tkhunny):

\(r = 2\sin(2\theta)\) \(r = 2(2\sin(\theta)\cos(\theta))\) \(r = 4\sin(\theta)\cos(\theta)\)

OpenStudy (tkhunny):

As in, I don't see that getting us anywhere. What can we do with mine?

OpenStudy (anonymous):

\[x^2 ++ y^2 = (\sin 2θ)^2 * (\cos θ)^2 + (\sin 2θ)^2 * (\sin θ_^2\] \[= (\sin 2θ)^2 = \sin 2θ * \sin 2θ = \sin 2θ (2 \sin θ \cos θ)\]

OpenStudy (tkhunny):

You're making a mess. Is that getting anywehre? \(r = 4\sin(\theta)\cos(\theta)\) If \(r \ne 0\) \(r\cdot r^{2} = 4r\sin(\theta)r\cos(\theta)\) \(r\cdot (x^{2}+y^{2}) = 4yx\) Are we closer or still just wandering around?

OpenStudy (anonymous):

Still a bit confused

OpenStudy (tkhunny):

Remember the "Fundamental Clues". I didn't mention \(r^{2} = x^{2} + y^{2}\). These are our ONLY conversion tools. We have to manipulate things to get those structures. That's all I did. After the double angle formula, I just used the three Fundamental Clues.

OpenStudy (anonymous):

ahhh I seee it

OpenStudy (anonymous):

Thank you @tkhunny

OpenStudy (tkhunny):

You still have to get rid of the last 'r'. \(r = \sqrt{x^{2} + y^{2}}\). You may wish to square some things and make it look prettier.